Database last updated on Mon, 02 Feb 2026 01:10:59 +0000 by localhost

Instrument constant changes for the week before 2024047 (4 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 14 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
08520240479613.7

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
0852024047121.4

Non-critical warnings from b-files' comments for days between 2024047 and 2024047 (0 sec, 6 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
hg:DEL key pressed399B04724.085TORON#08513:30:071-1
gf:HOME key pressed299B04724.085TORON#08515:26:161-1
pd:Going to menu299B04724.085TORON#08515:26:181-1
dsp:Doing dispersion test with Hg (UV port)299B04724.085TORON#08517:02:101-1
checkhg:Last HG had to adjust the micrometer by at least 1299B04724.085TORON#08518:11:371-1
fr:Micrometer #1 reference found at step 8758 (+CAL)= 9782 Now set to o3 offset = 8440199B04724.085TORON#08520:17:50978210009000

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssl
047TORONTO#085menu,Waiting until ZA = 170 now 138.1,Waiting until ZA = 170 now 139.5,Waiting until ZA = 170 now 140.8,Waiting until ZA = 170 now 142,Waiting until ZA = 170 now 143.1,Waiting until ZA = 170 now 144.2,Waiting until ZA = 170 now 145.2,Waiting until ZA = 170 now 146,Waiting until ZA = 170 now 146.7,Waiting until ZA = 170 now 147.3,Waiting until ZA = 170 now 147.8,Waiting until ZA = 170 now 148.1,Waiting until ZA = 170 now 148.312492611
(0 sec)

Dead time (0 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B04724.0851#0852831